A company issued a dividend of $1.00 (D0) this year, which is expected to grow at 15% per year for the next 2 years and 5% per year thereafter. The required rate of return is 15% per year. Use the two-growth dividend discount model to complete the following table.

Item Value
N   
rs   %
g1   %
g2   %
D0 $1.00
D1 $   
D2 $   
D3 $   
P0 $   
Answer #1

Given information:

rs = 15%

g1 = 15%

g2 = 5%

D0 = $1.00

D1 = D0 * ( 1 + g1) = 1.00 * (1+0.15) = $1.15

D2 = D1 * ( 1 + g1) = 1.15 * ( 1+0.15) = $1.32

D3 = D2 * ( 1 + g2) = 1.32 * (1+0.05) = $1.39 (The dividend growth rate changes in the third year, hence g2 is used instead of g1)

P0 = PV of dividends paid for next 2 years + PV of stock calculated using Gordon's constant growth model at the end of year 2.

= D1 / (1+rs) + D2 / (1+rs)2 + [ D3 / (r-g2) ] / (1+rs)2

= 1.15 / (1+0.15) + 1.32 / (1+0.15)2 + [ 1.39 / (0.15-0.05) ] / (1+0.15)2

= $12.50

The completed table is given below

Item Value
N infinity
rs 15%
g1 15%
g2 5%
D0 $1.00
D1 $1.15
D2 $1.32
D3 $1.39
P0 $12.50